# DISCOVER: Data-driven Identification of Sub-activities via Clustering and Visualization for Enhanced Recognition of Human Activities in Smart Homes

### v1

@@ 7.2. Future Work
While our approach demonstrates promising results, it has certain limitations which can be addressed in future work. First, our current method processes simple sensor activation sequences without considering additional context such as sensor location, time of day, or activation duration. Future work could enrich these input sequences by incorporating spatial and temporal contextual information using TDOST-style sentence embeddings (Thukral et al., 2024), which could lead to an improved overall performance. Second, the use of fixed sliding windows (i.e., 20 sensor readings) limits contextual understanding, as the model does not account for sensor events immediately preceding or following the window. To address this limitation for model training, future work could either increase the window size or append pre-text and post-text windows to provide additional temporal context. Additionally, integrating this extended context into our visualization tool would help annotators better interpret the sensor sequences. Third, our current approach requires manually creating a hierarchical label tree for each dataset, a process that is both labor-intensive and dataset-specific. While we want to preserve researchers’ flexibility in choosing the level and granularity of activities, our method could benefit from a unified hierarchical taxonomy of activities, which could be learned from the data by processing multiple heterogeneous datasets all at once. Fourth, our current manual labeling process assumes that annotators must review every sample and select a label from an extensive list of possibilities, which is both time-consuming and prone to inconsistencies. To address this, we could partially automate the process by leveraging large language models to pre-populate candidate labels. For example, by encoding sensor sequences as descriptive, TDOST-style sentences (Thukral et al., 2024), we could automatically summarize cluster centroids, thereby presenting annotators with pre-generated labels for validation rather than requiring them to select the labels from the full list. Finally, the approach assumes that activity patterns remain stable over time, while in real life significant changes in resident behavior may require periodic re-training or re-clustering. Addressing these limitations in future work will further enhance the adaptability and robustness of our framework for real-world smart home environments.
@@ 8. Summary and Conclusion
In this paper, we introduced DISCOVER, a method for discovering fine-grained sub-activities from unlabeled smart home sensor data without relying on pre-segmentation. Our pipeline is organized into two core steps: Clustering and Labeling. The Clustering step consists of:
  • Encoder Pre-Training: We leverage a pre-trained BERT model adapted with sensor-specific tokens and train it using a masked language modeling (MLM) objective to generate context-rich embeddings for raw sensor sequences.
  • Clustering Model Fine-Tuning: Using the SCAN loss function, we fine-tune these embeddings to form more homogeneous and distinct clusters of sensor sequences.
The Labeling step comprises:
  • Cluster Centroid Annotation: Representative sequences from each cluster are visualized with a custom tool, enabling expert annotators to assign meaningful sub-activity labels to the centroids.
  • Label Propagation: The centroid labels are propagated to all sequences within their respective clusters, resulting in a fully labeled dataset with minimal manual effort.
  • Re-annotation of Original Time-Series Data: Finally, these propagated labels are mapped back onto the original time-series data, preserving temporal continuity and facilitating the analysis of longitudinal activity patterns.
Our approach addresses important challenges in HAR, including the high cost and effort of manual data annotation, the limitations of coarse activity labels, and the need for scalable and generalizable models. DISCOVER offers an open source tool that facilitates the HAR annotation and re-annotation process and enables the dynamic discovery and validation of sub-activities, thus capturing a broader spectrum of behaviors observed in real homes.
@@ Appendix A Appendix
Following Section 5.2.3, we present the tSNE projections of DISCOVER embeddings for all three datasets: Milan on Fig. 7, Aruba on Fig. 8, and Cairo on Fig. 9. The left panel shows the embeddings colored by CASAS label; the right one shows the same exact embeddings colored by SCAN cluster labels from 0 to 19. Note how the Other label (gray color) is mostly scattered around . Also note that Cairo, which has 81% of its data points labeled as Other (Table 1), has most of its clusters fully gray.
@@
A.3. DISCOVER Sub-Activity Labels
Fig. 11 shows the list of all DISCOVER sub-activity labels that we provided the annotators with. The activities are organized hierarchically by categories, e.g. Single Room Activity -¿ Kitchen Activity -¿ Movement in Kitchen -¿ Movement Near Stove. These labels were used to populate the labeling drop-down menu in DISCOVER visualization tool shown in Fig. 3.
Figure 11. DISCOVER sub-activity hierarchy. This list of manually-curated labels was used to populate the labeling drop-down menu in DISCOVER visualization and labeling tool.
A.4. Individual Cluster Sub-Activity Labels
@@
Figure 12. Sub-activity labels assigned to each cluster manually labeled clusters for all households. votes column shows how many annotations aligned with this particular label; level up column provides a parent label.
@@ A.5. Household Layouts
To accurately visualize and analyze sensor data, we utilized the household layouts from (Hiremath et al., 2022; Thukral et al., 2024; Cook et al., 2009). We re-rendered these layouts to simplify the visuals and ensure compatibility with our DISCOVER visualization tool.

### v3

@@ 7.4. Future Work
While our approach demonstrates promising results, it has certain limitations which can be addressed in future work. First, our current method processes simple sensor activation sequences without considering additional context such as sensor location, time of day, day of the week, or activation duration. Future work could enrich these input sequences by incorporating spatial and temporal contextual information using either special tokens or descriptive embeddings. This integration could facilitate the discovery of more nuanced behavioral patterns, such as differentiating between daytime and nocturnal bathroom visits or distinguishing unique weekday versus weekend routines.
Second, while the current framework successfully identifies Patterns of Daily Living (PDL), the organization of these patterns into a coherent hierarchy still relies on domain-specific manual mapping. Future work should focus on moving beyond fixed label trees toward the automated discovery of data-driven taxonomies. By processing multiple heterogeneous datasets simultaneously, DISCOVER could potentially learn a universal structural representation of human behavior that transcends individual household variations. Such a transition from discrete pattern recognition to a systematic, hierarchical taxonomy would provide a deeper, data-grounded understanding of the fundamental building blocks of human routines. This evolution would enable the longitudinal analysis of behavior not just as a sequence of activities, but as a dynamic interplay of multi-level behavioral motifs, ultimately allowing for more robust cross-environment generalization.
Finally, while our current approach reduces the volume of data requiring manual annotation, the selection of semantic labels for each discovered cluster still imposes a cognitive burden on the expert. To further streamline this process, future work could explore the alignment of learned sensor embeddings with natural language representations in a shared latent space. Rather than utilizing Large Language Models (LLMs) as black-box annotators (an approach vulnerable to factual hallucinations and a lack of environmental grounding (Huang et al., 2025)), we propose leveraging LLMs to rank and surface candidate descriptors derived from the latent spatial and temporal regularities identified by DISCOVER. By surfacing the most probable behavioral motifs for a given cluster, we shift the human role from exhaustive classification to an efficient verification process. This data-grounded alignment would maintain high annotation fidelity while significantly reducing the time and expertise required to label complex, longitudinal sensor streams across diverse households.
@@ 8. Summary and Conclusion
In this paper, we introduced DISCOVER, a framework for discovering Patterns of Daily Living from continuous ambient-sensor streams without any supervision signal. By combining self-supervised Transformer pre-training with representation-aware clustering (SCAN) and a lightweight human-in-the-loop labeling workflow supported by an interactive visualization tool, DISCOVER discovers interpretable, household-specific Patterns of Daily Living (PDL) with minimal annotation effort, providing a data-grounded foundation for longitudinal behavioral analysis. Across three CASAS homes, we show that the resulting clusters are coherent and homogeneous (high inter-rater and cluster agreement), uncover fine-grained structure beyond predefined ADL vocabularies, and achieve classification performance comparable to fully supervised baselines while using orders of magnitude fewer labels.
Together, these results position DISCOVER as a robust complement to traditional HAR systems, providing the granular resolution necessary to move beyond broad activity labels. By grounding behavioral patterns in a resident’s specific environment, our framework enables the detection of subtle, within-person habitual shifts, that serve as essential digital biomarkers for early-stage cognitive decline and longitudinal health monitoring.
@@ Appendix A Appendix
Following Section 5.4, we present the tSNE projections of DISCOVER embeddings for all three datasets: Milan on Fig. 7, Aruba on Fig. 8, and Cairo on Fig. 9. The left panel shows the embeddings colored by CASAS label; the right one shows the same exact embeddings colored by SCAN cluster labels from 0 to 19. Note how the Other label (gray color) is mostly scattered around. Also note that Cairo, which has 81% of its data points labeled as Other (Table 1), has most of its clusters fully gray.
@@
A.3. Patterns of Daily Living: Labels
Fig. 11 shows the list of all DISCOVER PDL labels that we provided the annotators with. The activities are organized hierarchically by categories, e.g. Single Room Activity -¿ Kitchen Activity -¿ Movement in Kitchen -¿ Movement Near Stove. These labels were used to populate the labeling drop-down menu in DISCOVER visualization tool shown in Fig. 3.
Figure 11. DISCOVER Patterns of Daily Living hierarchy. This list of manually-curated labels was used to populate the labeling drop-down menu in DISCOVER visualization and labeling tool.
A.4. Individual Clusters
@@
Figure 12. Patterns of Daily Living labels assigned to each cluster manually labeled clusters for all households. votes column shows how many annotations aligned with this particular label; level up column provides a parent label.
@@ A.5. Household Layouts
We re-rendered the original CASAS (Cook et al., 2009) household layouts to simplify them visually and make certain household features like furniture stand out more prominently, following the insights from (Hiremath et al., 2022).  Fig. 13 shows the simplified layouts for Milan, Aruba, and Cairo.